### v1

# P[$\arraybackslashp#1$

Abstract
Large Language Models (LLMs) have rapidly advanced and are widely adopted across diverse fields. Due to the substantial computational cost and data requirements of training from scratch, many developers choose to fine-tune or modify existing open-source models. While most adhere to open-source licenses, some falsely claim original training despite clear derivation from public models. This raises pressing concerns about intellectual property protection and highlights the need for reliable methods to verify model provenance. In this paper, we propose GhostSpec, a lightweight yet effective method for verifying LLM lineage without access to training data or modification of model behavior. Our approach constructs compact and robust fingerprints by applying singular value decomposition (SVD) to invariant products of internal attention weight matrices, effectively capturing the structural identity of a model. Unlike watermarking or output-based methods, GhostSpec is fully data-free, non-invasive, and computationally efficient. It demonstrates strong robustness to sequential fine-tuning, pruning, block expansion, and even adversarial transformations. Extensive experiments show that GhostSpec can reliably trace the lineage of transformed models with minimal overhead. By offering a practical solution for model verification and reuse tracking, our method contributes to the protection of intellectual property and fosters a transparent, trustworthy ecosystem for large-scale language models.
1 Introduction
Large Language Models (LLMs) have rapidly advanced and become foundational infrastructure for a wide range of applications across research and industry (achiam2023gpt; yang2025qwen3; dubey2024llama). However, the high computational and data costs of training LLMs from scratch (workshop2022bloom) have led many developers to build their own models by modifying existing open-source LLMs through techniques such as instruction fine-tuning, continued pre-training, model merging, and compression (yang2024model; zhu2024survey; tang2025spindlekv). While most developers comply with open-source licenses, there have been instances of falsely claiming to have trained models “from scratch” when they are in fact repackaged or fine-tuned versions of public models (e.g., Llama3-V and MiniCPM-Llama3-V 2.5) (yao2024minicpm). This trend raises serious concerns regarding model plagiarism and intellectual property violations, highlighting the urgent need for reliable tools to verify model lineage.
To address these concerns, researchers have proposed various model identification methods (sun2023deep), which can be broadly classified into black-box and white-box approaches. Black-box methods identify models without accessing their weights, using techniques like behavioral fingerprinting and watermarking. However, these approaches are often sensitive to randomness, adversarial changes, or require intrusive pipeline modifications. In contrast, white-box methods leverage internal parameters. While representation-based techniques analyze hidden states or gradients, they depend on data access and are computationally expensive. Direct weight comparisons are simpler but fragile under fine-tuning or pruning.
In this work, we propose GhostSpec, a simple, data-free, and robust white-box method for verifying the lineage of LLMs. Our key insight is that the spectral structure of weight matrices encodes intrinsic information about a model’s origin, remaining stable under various modifications. Specifically, we apply singular value decomposition (SVD) to internal matrix products within the attention mechanism, including the query-key and value-output weight products. From the resulting spectra, we select the most significant singular values based on effective rank to compute similarity. To handle architectural variations in depth, such as those resulting from layer pruning or expansion, we develop the Penalty-based Optimal Spectral Alignment (POSA) algorithm that finds the best layer-wise correspondence between models. This yields a quantitative similarity score robust to differences in depth and architectural variations. Unlike black-box or representation-based white-box methods, GhostSpec is data-independent, requires no model modification, and has minimal computational cost.
Figure 1: GhostSpec extracts singular value spectra from each layer’s attention products to form spectral fingerprints. A pairwise MSE distance matrix is computed, and a penalty-based alignment algorithm matches layers across models of different depths. The final similarity score distinguishes between related and independently trained models.
Our contributions are summarized as follows:
  • We propose GhostSpec, a lightweight white-box method for verifying LLM lineage directly from model weights. It requires no training data or architectural changes, and remains robust to fine-tuning and pruning—offering a practical solution for provenance verification and IP protection in open-source LLMs.
  • We introduce spectral fingerprints based on invariant matrix products within the attention mechanism, which are intrinsically robust to functionality-preserving scaling and permutation transformations. Additionally, we develop the Penalty-based Optimal Spectral Alignment (POSA) algorithm to enable robust comparison across models with varying depths and architectures.
  • We conduct extensive experiments demonstrating that GhostSpec effectively distinguishes derivative models from independently trained ones, even under challenging modification scenarios, validating its utility and reliability.
2 Related Work
Efforts to verify model lineage can be broadly classified into black-box and white-box approaches, depending on whether internal model access is required.
2.1 Black-Box Identification
Black-box methods operate without access to model weights and are suitable for closed-source, API-only models. They include behavioral fingerprinting and watermarking.
Behavioral Fingerprinting.
These passive methods identify model signatures from natural outputs. Approaches analyze stylistic or statistical patterns in generated text, or use crafted prompts to probe responses (pasquini2024llmmap; mcgovern2024your; sam2025predicting). Some rely on output logits or top-k probabilities to define a unique model space (yang2024fingerprint). However, such methods are sensitive to decoding randomness and vulnerable to adversarial paraphrasing.
Watermarking.
Watermarking embeds a detectable signal in model outputs, either via instruction tuning (xu2024instructional) or token-level perturbation (kirchenbauer2023watermark; nagatsuka2025nested). These signals can be verified statistically, but require the model creator’s cooperation and can be invalidated by output editing or algorithm exposure.
2.2 White-Box Identification
White-box approaches leverage internal model weights or activations, typically by computing similarity between weights, intermediate representations, or gradients.
Representation-based Fingerprinting.
These methods analyze hidden representations and generally require input data. Techniques such as CKA similarity or gradient statistics have been used to reveal shared training origins (zhang2024reef; wu2025gradient; liang2025origin). While effective, these methods are computationally intensive, data-dependent, and may raise concerns regarding potential correlations with training data.
Weight-based Fingerprinting.
This line of work focuses on static, data-free analysis of model weights. Prior methods include visualizing invariant structural features (zeng2024huref) or analyzing layer-wise statistics (yoon2025intrinsic). Our proposed GhostSpec method falls into this category, capturing deeper structural information by leveraging the full singular value spectrum of invariant matrix products. The Penalty-based Optimal Spectral Alignment(POSA) algorithm is further introduced to robustly trace model ancestry under various transformations.
Figure 2: Average MSE of normalized singular values from Q/K/V/O projections. The spectral distance from Llama-2-7b to its fine-tuned variants (blue) is negligible, while the distance to unrelated models (red) is large, confirming the fingerprint’s robustness against fine-tuning.
Figure 3: The layer-wise trend of the mean of normalized singular values for various models. Models with a shared lineage (e.g., Llama-2-7b-hf and its variants) exhibit highly correlated trends, while unrelated models show divergent patterns.
3 Preliminaries
staats2024small employ Random Matrix Theory (RMT) to analyze the singular value spectrum of weight matrices in pre-trained language models. Their analysis reveals significant deviations from the Marchenko–Pastur distribution expected from randomly initialized weight matrices, especially in the region of large singular values. These outliers correspond to dominant directions in the weight space and are strongly associated with the model’s learned representations. Removing them leads to a substantial increase in perplexity, indicating that they form a stable backbone of the model’s identity.
Fine-tuning, in contrast to training from scratch, refines rather than rebuilds the model’s internal structure. staats2024small demonstrate that fine-tuning primarily affects directions associated with small singular values. A key asymmetry is observed: pruning small singular values after fine-tuning leads to significantly greater performance degradation than pruning them before fine-tuning. This indicates that fine-tuning updates are concentrated in low-magnitude spectral components.
Together, these findings suggest spectral stability: large singular values, encoding foundational pre-trained knowledge, remain stable throughout adaptation and anchor the model’s global behavior. In contrast, fine-tuning introduces targeted, low-rank modifications reflected in small singular values. Based on this observation, we hypothesize that the large singular value spectrum of attention matrices serves as a robust fingerprint of a model, largely invariant to fine-tuning.
3.1 Experimental Design
To empirically test this hypothesis, we designed an experiment centered on Llama-2-7b-hf. We compared this primary model against two distinct groups:
  • Fine-tuned Variants: A group of known direct descendants of Llama-2-7b (touvron2023llama), including Llama-2-7b-chat-hf, Vicuna-7b-v1.5, Llemma_7b (azerbayev2023llemma), and Xwin-LM-7B-V0.2 (xwin-lm).
  • Unrelated Models: A control group of architecturally distinct models, including Mistral-7B-v0.1 (DBLP:journals/corr/abs-2310-06825), Pythia-6.9b (biderman2023pythia), Yi-1.5-6B (young2024yi), Internlm2-base-7b (cai2024internlm2).
Our objective is to quantify the spectral distance between the base model and each model in these groups, expecting minimal distance for fine-tuned variants and a significantly larger distance for unrelated models.
3.2 Quantifying Spectral Similarity
To quantify the similarity between the singular value spectra of two models $A$ and $B$, we define a compact layer-wise distance metric.
For each projection type $p\in\{q,k,v,o\}$ and each layer $i\in\{1,\dots,L\}$, we extract the singular value vectors $\mathbf{s}_{p,A}^{(i)}$ and $\mathbf{s}_{p,B}^{(i)}$ from the corresponding weight matrices. We first truncate both vectors to their minimum effective rank $r_{p}^{(i)}$ and then apply min-max normalization to map them into $[0,1]$. The spectral distance at layer $i$ and projection $p$ is then defined as the mean squared error (MSE) between the normalized, truncated vectors:
$\begin{split}d_{p}^{(i)}(A,B)=\frac{1}{r_{p}^{(i)}}\left\lVert\right.&\text{norm}(\text{trunc}(\mathbf{s}_{p,A}^{(i)}))\\ &\left.-\text{norm}(\text{trunc}(\mathbf{s}_{p,B}^{(i)}))\right\rVert_{2}^{2},\end{split}$ (1)
where $\text{trunc}(\cdot)$ denotes truncation to rank $r_{p}^{(i)}$ and $\text{norm}(\cdot)$ denotes min-max normalization.
The overall spectral distance is computed by averaging over all layers and projection types:
$D(A,B)=\frac{1}{4L}\sum_{p\in\{q,k,v,o\}}\sum_{i=1}^{L}d_{p}^{(i)}(A,B).$ (2)
3.3 Empirical Validation
As shown in Figure 2, fine-tuned models exhibit a low average spectral MSE from their base models, indicating near-identical structural characteristics. In contrast, unrelated models yield significantly higher distances, highlighting their structural dissimilarity.
These results validate that the singular value spectrum is a stable and intrinsic property of a model, largely preserved through fine-tuning. This motivates the design of a robust fingerprinting framework capable of verifying model lineage under various transformations.
[…]
Vicuna-7b-v1.5
Llemma_7b
Llama-2-7b
-scaled
Llama-2-7b
-permuted
Pruned-50%
-Retrained
Pruned-70%
-Retrained
[…]
Sheared-Llama
1.3B
Sheared-Llama
2.7B
Llama2-7b-func
-call-slerp
Camelidae-8x7B
Qwen2.5-7B
OPT-6.7b
[…]
OpenHermes-2.5
-Mistral-7B
Triunvirato-7b
Chunky-Lemon
-Cookie-11B
OpenHermes-2.5
-Mistral-7B-pruned50
Qwen2.5-7B Yi-1.5-6B
[…]
Table 1: Comprehensive comparison of fingerprinting methods against various derivative and unrelated models, with Llama-2-7b and Mistral-7B as primary models. The table evaluates robustness to fine-tuning, architectural dissimilarity, compression, merging, expansion, and adversarial transformations. Similarity scores are color-coded based on method-specific thresholds: $x$ indicates a score above the threshold (positive classification), while $x$ indicates a score below it (negative classification).
Algorithm 1 Penalty-based Optimal Spectral Alignment
Input: Distance matrix $D\in\mathbb{R}^{N\times M}$ (assume $N\leq M$), gap penalty $\rho$.
Output: Average distance along the optimal path.
1: Initialize cost matrix $C\in\mathbb{R}^{N\times M}$ and backtrack matrix $B\in\mathbb{Z}^{N\times M}$.
2:for $j=0$ to M1M-1 do
3:  $C[0,j]\leftarrow D[0,j]$ {Base case for the first layer.}
4:end for
5:for $i=1$ to N1N-1 do
6:  for $j=i$ to M1M-1 do
7:   Find $k^{*}=\arg\min\limits_{i-1\leq k<j}\left(C[i-1,k]+(j-k-1)\rho\right)$
8:   $C[i,j]\leftarrow C[i-1,k^{*}]+(j-k^{*}-1)\rho+D[i,j]$
9:   $B[i,j]\leftarrow k^{*}$
10:  end for
11:end for
[…]
13: Reconstruct optimal path $P$ by backtracking from $(N-1,j_{\text{end}})$ using $B$.
14:return $\frac{1}{|P|}\sum\limits_{(i,j)\in P}D[i,j]$ {Avg. MSE on path.}
4 Methodology
In practice, model weights may undergo transformation attacks such as permutation or scaling, which preserve the model’s functionality while significantly altering the raw weight distribution (zhang2024reef). Direct comparison of the singular values of individual attention matrices $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ is therefore highly vulnerable to such obfuscation techniques. To address this, we propose a robust spectral fingerprinting methodology built on two core principles: (1) an invariant fingerprint derived from composite matrix products, and (2) two complementary similarity metrics, Ghost-mse and Ghost-corr, designed to jointly capture fine-grained variations and macroscopic structural properties.
4.1 The GhostSpec Fingerprint
To quantify and compare the structural properties of Transformer-based language models, we first define an invariant spectral fingerprint that is robust to permutation and scaling transformations. For a given model $M$ with $L$ layers, we focus on the attention-related weight matrices: $W_{q}^{(i)}$, $W_{k}^{(i)}$, $W_{v}^{(i)}$, and $W_{o}^{(i)}$ from each layer $i\in\{1,\dots,L\}$.
We define two invariant matrices per layer whose singular value spectra are resilient to adversarial attacks (see Appendix A for proof):
$M_{qk}^{(i)}=W_{q}^{(i)}(W_{k}^{(i)})^{T}\quad\text{and}\quad M_{vo}^{(i)}=W_{v}^{(i)}W_{o}^{(i)}.$ (3)
The fingerprint for each layer is composed of the two singular value vectors derived from these products. The layer fingerprint $\mathcal{S}_{M}^{(i)}$ is defined as the tuple:
$\mathcal{S}_{M}^{(i)}=\left(\mathbf{s}_{qk,M}^{(i)},\;\mathbf{s}_{vo,M}^{(i)}\right),$ (4)
where $\mathbf{s}_{p,M}^{(i)}=\mathrm{SVD}(M_{p,M}^{(i)})$ for $p\in\{qk,vo\}$. The complete model fingerprint $\mathcal{F}_{M}$ is the sequence of these layer fingerprints across all layers.
4.2 Similarity Metrics
Given the GhostSpec fingerprints of two models, $A$ and $B$, we introduce two complementary metrics to quantify their similarity.
Fine-grained Similarity: GhostSpec-mse.
This metric performs a detailed, layer-by-layer comparison of the singular value vectors to populate an aggregate distance matrix $D_{\text{avg}}\in\mathbb{R}^{N\times M}$. Each entry $(D_{\text{avg}})_{ij}$ represents the average spectral distance between layer $i$ of model A and layer $j$ of model B, computed as the mean of the Mean Squared Error (MSE) over the invariant components:
$(D_{\text{avg}})_{ij}=\frac{1}{2}\sum_{p\in\{qk,vo\}}\frac{1}{r_{p,ij}}\left\|\hat{\mathbf{s}}_{p,A}^{(i)}-\hat{\mathbf{s}}_{p,B}^{(j)}\right\|_{2}^{2}$ (5)
where:
  • $p$ is the invariant product type ($qk$ or $vo$).
  • $r_{p,ij}=\min(\text{eff\_rank}(\mathbf{s}_{p,A}^{(i)}),\text{eff\_rank}(\mathbf{s}_{p,B}^{(j)}))$ is the minimum of the effective ranks of the two singular value vectors being compared.
  • $\hat{\mathbf{s}}$ denotes a processed singular value vector. The processing involves first truncating the original vector $\mathbf{s}$ to its top $r_{p,ij}$ values, and then applying min-max normalization to scale the result to the $[0,1]$ range.
To handle models with different depths ($N\neq M$), we apply our Penalty-based Optimal Spectral Alignment (POSA) algorithm 1 to find the minimum-cost alignment path through $D_{\text{avg}}$. The raw similarity score, $d_{\text{path}}$, is the average MSE along this optimal path. Finally, we convert this distance into a normalized similarity score using an inverted Sigmoid transformation:
$\mathrm{Sim}_{\text{MSE}}(A,B)=1-\frac{1}{1+e^{-k(d_{\text{path}}-\tau)}},$ (6)
where $\tau$ is an empirical discrimination threshold and $k$ is a steepness factor. A score approaching 1.0 indicates high similarity.
Lightweight Similarity: GhostSpec-corr.
This metric captures the macroscopic relationship between models by comparing the overall trend of their spectral properties. For each layer $i$ and component $p$, we compute a single scalar value, $\mu_{p,M}^{(i)}$, which is the mean of the top-$K$ normalized singular values.
This process yields two trend sequences for each model: $\boldsymbol{\mu}_{qk,M}$ and $\boldsymbol{\mu}_{vo,M}$. After aligning the sequences to handle differing depths, we concatenate them and compute the final similarity score using the distance correlation coefficient:
$\begin{split}\mathrm{Sim}_{\text{Corr}}(A,B)=\mathrm{dCor}\big(&[\boldsymbol{\mu}_{qk,A};\boldsymbol{\mu}_{vo,A}],\\ &[\boldsymbol{\mu}_{qk,B};\boldsymbol{\mu}_{vo,B}]\big).\end{split}$ (7)
This score measures the correlation of the models’ high-level spectral evolution, providing a computationally efficient indicator of shared lineage.
Figure 4: Maximum F1 scores for each method on our 63-pair dataset. Both GhostSpec variants clearly outperform all baseline methods in accurately distinguishing related from unrelated models.
5 Experiments
To evaluate the effectiveness and robustness of our proposed GhostSpec, we conduct a comprehensive suite of experiments designed to emulate real-world scenarios of model reuse, modification, and transformation.
5.1 Experimental Setup
Dataset Construction.
We constructed a comprehensive dataset consisting of 63 model pairs, using Llama-2-7b and Mistral-7B as the primary base models. This dataset covers a wide range of transformations, including fine-tuning, compression, merging, and expansion, with ground-truth labels indicating whether each pair is related or unrelated.
Evaluation Protocol.
To ensure a fair and rigorous comparison with baseline methods, we establish a standardized evaluation protocol. For each method, a model pair is classified as ’related’ if its similarity score exceeds a certain threshold. We determine the optimal threshold $\tau^{*}$ for each method by finding the value that maximizes the F1-score against the ground truth labels (1 for related, 0 for unrelated):
$\tau^{*}=\operatorname*{arg,max}_{\tau\in\mathcal{S}}\text{F1}(\hat{y}_{i}(\tau),y_{i}).$ (8)
This protocol allows us to evaluate each method at its peak performance, ensuring a fair comparison.
Baseline Methods.
We compare GhostSpec against representative fingerprinting methods across multiple paradigms.
  • Data-Aware Baselines: These methods require input data to generate outputs or internal representations for analysis. This category includes black-box approaches like QueRE (sam2025predicting) and Logits (yang2024fingerprint), which analyze model outputs, as well as white-box methods like REEF (zhang2024reef), which measures hidden representation similarity. A key characteristic of these methods is their reliance on data, which increases computational cost and necessitates curated datasets.
  • Data-Free Baselines: These methods operate directly on static model weights without requiring any input data. This category includes methods like PCS (zeng2024huref), which analyzes invariant submatrices within transformer layers.
5.2 Main Results
The overall classification performance of each method, measured by its maximum F1-score on our dataset, is summarized in Figure 4. Our two proposed variants, GhostSpec-corr (F1 = 0.9730) and GhostSpec-mse (F1 = 0.9867), achieve a clear lead, outperforming all data-aware and data-free baselines. Table  1 presents a selection of the detailed similarity scores, color-coded using these optimal thresholds to visualize the classification results. The following sections detail the performance across specific modification scenarios.
Robustness to Fine-Tuning.
Fine-tuning is a common source of model derivation. We evaluate similarity between Llama-2-7B and several of its fine-tuned variants: Llama-2-7B-Chat-HF, Vicuna-7B-v1.5, Xwin-LM-7B-V0.2, and Llemma-7B, etc. Despite differences in training data and objectives, GhostSpec consistently produces high similarity scores.
Robustness to Model Pruning.
We assess GhostSpec’s robustness under model pruning. For structured pruning, we consider methods that remove layers based on importance, such as ShortGPT (men2024shortgpt) and Sheared Llama (xia2023sheared). For unstructured pruning, we evaluate representative importance-based methods that prune weights according to their significance, including Wanda (sun2023simple) and SparseGPT (frantar2023sparsegpt). Prior work (ma2025model) identifies critical sparsity thresholds near 35% (structured) and 60% (unstructured), beyond which model behavior and redundancy shift significantly. We thus evaluate sparsity levels at 30%, 50%, and 70%, finding that GhostSpec consistently and accurately recovers the source model across all settings.
Robustness to Model Merging.
Model merging combines weights from multiple source models to create a new model. We evaluate GhostSpec on various merging approaches, including direct weight averaging, methods that reduce parameter interference through pruning and sign conflict resolution (yadav2023ties), as well as techniques that merge models by fusing weights or aligning output distributions (goddard-etal-2024-arcees; wan2024knowledge). Across all these methods, GhostSpec consistently detects strong similarity between the merged models and their original parents.
Robustness to Model Upcycling.
Model upcycling expands model capacity by introducing new components, such as additional layers or sparsely activated expert modules, including techniques that convert dense LLMs into sparse MoEs through Parameter-Efficient Sparsity Crafting (PESC) (wu2024parameter). Despite architectural modifications, GhostSpec consistently yields high similarity scores with the original base models.
Robustness to Permutation and Scaling Transformations.
To evaluate invariance under adversarial weight manipulations, we apply random permutations to the hidden dimensions of MLP weight matrices and introduce random scaling factors to the attention layers. Despite these aggressive transformations, GhostSpec consistently recognizes the original and perturbed models as similar, validating its theoretical invariance guarantees.
Robustness to Mistral-based Derivatives.
To further validate the architectural generalizability of GhostSpec, we extend our evaluation to the Mistral model family beyond the Llama series. This evaluation covers a diverse set of Mistral-based derivatives, including instruction-tuned models, DPO-optimized variants (nous2024hermes), compressed versions using pruning, or sparsification, as well as models with structural changes such as merging, interpolation, and MoE-style expansion. Despite the diversity of these modifications, GhostSpec consistently maintains accurate and reliable lineage detection across all cases, demonstrating its strong robustness and cross-architecture transferability.
Discrimination Against Unrelated Models.
To evaluate the specificity of GhostSpec and minimize false positives, we tested its performance on architecturally unrelated models. We computed similarity scores between our base models (Llama-2-7b and Mistral-7B) and a diverse set of independently trained LLMs, including pythia-6.9b, opt-6.7b, among others(MosaicML2023Introducing; guo2024deepseek; zhang2022opt). GhostSpec consistently assigned near-zero similarity scores, reliably distinguishing true derivatives from unrelated models and demonstrating strong discriminative capability.
Figure 5: Pairwise structural similarity matrix of prominent open-source models computed using GhostSpec-mse. Higher scores indicate greater similarity.
Figure 6: MLP-based spectral fingerprint similarity between Llama-2-7b and various models using GhostSpec-mse, computed from singular values of MLP up_proj and down_proj weights. Higher scores indicate greater similarity.
5.3 Further Discussion
Resilience to Evasion Attacks.
We investigate an evasion strategy wherein an adversary fine-tunes a model with a custom loss to obscure its structural fingerprint. The objective is to maximize spectral distance from a victim model while preserving task performance.
To simulate this, we define a total loss combining task loss and a spectral divergence term:
$\mathcal{L}_{\text{total}}(\theta)={}$ $\mathcal{L}_{\text{task}}(\theta)-\lambda\cdot D_{\text{SVD}}(\mathcal{M}_{\theta},\mathcal{M}_{\text{victim}})$ (9)
where $D_{\text{SVD}}$ denotes the MSE between singular value spectra. The hyperparameter $\lambda$ balances task fidelity and spectral divergence.
We apply this method to adversarially fine-tune a Llama-3.2-1B-instruct model. Results show that it is difficult to significantly alter the singular value spectrum without hurting model performance.
These findings indicate that spectral features are intrinsically tied to model functionality, limiting the practicality of evasion via spectral manipulation.
Case Study: A High-Profile Lineage Dispute
We apply GhostSpec using Ghost-mse similarity to investigate the recently debated lineage of Pangu-Pro-MoE by comparing it with several open-source models, including the Qwen series and Llama variants.
As shown in Figure 5, GhostSpec finds that Pangu-Pro-MoE has the highest similarity with OpenMath-Nemotron-14B (a fine-tuned variant of Qwen2.5-14B) and Qwen2.5-14B, while showing negligible similarity to unrelated models such as Yi-1.5-9B and Llama-2-13b-hf.
These results indicate a potential lineage connection between Pangu-Pro-MoE and the Qwen2.5-14B family, though further evidence is needed for confirmation.
Analysis of MLP Module Spectra.
To examine whether spectral fingerprinting extends beyond attention layers, we analyze singular values of the MLP up-porj and down-proj weight matrices in each Transformer layer. We compute Ghost-mse similarity between Llama-2-7b-hf, its fine-tuned variants, and unrelated models.
As shown in Figure 6, MLP-based fingerprints effectively distinguish fine-tuned models (e.g., Vicuna-7b-v1.5, Llemma-7b) with similarity above 0.96 from unrelated models (e.g., Pythia-6.9b) showing near-zero similarity.
However, this approach incurs higher computational cost due to the larger size of MLP weight matrices and is less robust to common dense-to-MoE expansions where MLP layers are replaced by experts. Therefore, attention-based fingerprints remain more efficient and structurally stable for reliable lineage verification.
6 Conclusion
This paper tackles the challenge of verifying large language model lineage amid widespread undocumented reuse and potential plagiarism. We propose GhostSpec, a robust, data-independent, and efficient white-box method that constructs stable fingerprints from the singular value spectra of invariant attention matrix products, resilient to common model modifications and adversarial transformations. A penalty-based optimal path alignment algorithm enables handling architectural differences. Extensive experiments demonstrate that GhostSpec reliably identifies model ancestry across diverse transformations, including fine-tuning, pruning, merging, and architectural expansion. It accurately recovers known relationships and offers insights in complex cases. Moreover, the fingerprint’s tight coupling with model functionality makes evasion attempts impractical without degrading performance. Overall, GhostSpec offers a practical, trustworthy tool to protect intellectual property and improve transparency, advancing verifiable model provenance in open-source AI ecosystems.
Appendix A Theoretical Foundation of the GhostSpec Fingerprint
This appendix provides the detailed theoretical underpinning for the robustness of the GhostSpec fingerprint, as claimed in Section 4.1 of the main paper. We first provide a rigorous derivation of the functionality-preserving transformations our method is designed to be invariant to. We then provide a formal proof of this invariance, taking into account the practical convention of storing transposed weight matrices.
Throughout this section, we denote the conceptual weight matrices as $W_{q},W_{k},W_{v},W_{o}$. For computational efficiency (e.g., using row-major input vectors $x$), models typically store their transposed counterparts: $W_{q}^{\top},W_{k}^{\top},W_{v}^{\top},W_{o}^{\top}$. A projection is thus calculated as $x\cdot W^{\top}$. Our derivations will explicitly use this convention.
A.1 Functionality-Preserving Transformation Models
Our claim of robustness is based on the invariance of our fingerprint to transformations that modify model weights without altering the model’s input-output function. We now show in detail how these transformations preserve functionality.
FFN Permutation.
A standard FFN block with a gated linear unit (GLU) computes its output $y$ from an input $x$ as:
$y=\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot W_{\text{down}}^{\top},$
where $\odot$ is element-wise multiplication and $\sigma$ is an activation function.
An adversary can apply a permutation matrix $P$ to the intermediate neurons. The transformed (and stored) weights $\tilde{W}^{\top}$ are:
$\tilde{W}_{\text{up}}^{\top}=W_{\text{up}}^{\top}P,\quad\tilde{W}_{\text{gate}}^{\top}=W_{\text{gate}}^{\top}P,\quad\tilde{W}_{\text{down}}^{\top}=P^{-1}W_{\text{down}}^{\top}.$
The output with the transformed weights, $\tilde{y}$, is:
$\tilde{y}$ $=\left(x\cdot\tilde{W}_{\text{up}}^{\top}\odot\sigma(x\cdot\tilde{W}_{\text{gate}}^{\top})\right)\cdot\tilde{W}_{\text{down}}^{\top}$
$=\left(x\cdot W_{\text{up}}^{\top}P\odot\sigma(x\cdot W_{\text{gate}}^{\top}P)\right)\cdot(P^{-1}W_{\text{down}}^{\top})$
$=\left(\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot P\right)\cdot(P^{-1}W_{\text{down}}^{\top})$
$=\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot I\cdot W_{\text{down}}^{\top}=y.$
The output $y$ is identical, hence functionality is preserved.
V-O Per-Head Transformation.
The context vector $c$ from a single attention head is computed by projecting the value vectors $v$ with the output weight matrix $W_{o}$. The value vectors themselves are projections of the input $x$, i.e., $v=x\cdot W_{v}^{\top}$. Thus, the final output $y$ of the head is:
$y=(\text{AttentionScores}\cdot(x\cdot W_{v}^{\top}))\cdot W_{o}^{\top}.$
The transformation involves a block-diagonal invertible matrix $C$ applied to the conceptual weights as:
$\tilde{W}_{v}=CW_{v},\quad\tilde{W}_{o}=W_{o}C^{-1}.$
The corresponding stored weights are:
$\tilde{W}_{v}^{\top}=W_{v}^{\top}C^{\top},\quad\tilde{W}_{o}^{\top}=(C^{-1})^{\top}W_{o}^{\top}.$
The output becomes:
$\tilde{y}$ $=(\text{AttentionScores}\cdot(x\cdot\tilde{W}_{v}^{\top}))\cdot\tilde{W}_{o}^{\top}$
$=(\text{AttentionScores}\cdot(x\cdot W_{v}^{\top}C^{\top}))\cdot((C^{-1})^{\top}W_{o}^{\top})$
$=(\text{AttentionScores}\cdot x\cdot W_{v}^{\top})\cdot(C^{\top}(C^{-1})^{\top})\cdot W_{o}^{\top}$
$=(\text{AttentionScores}\cdot x\cdot W_{v}^{\top})\cdot I^{\top}\cdot W_{o}^{\top}=y.$
Q-K RoPE-Compatible Transformation.
Self-attention output is computed as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (10)
where $Q=xW_{q}^{\top}$ and $K=xW_{k}^{\top}$.
To construct a transformation that preserves this output while remaining compatible with RoPE, we introduce a doubly block-diagonal matrix $P$ and define:
$\tilde{W}_{q}=PW_{q},\quad\tilde{W}_{k}=(P^{\top})^{-1}W_{k}.$ (11)
The corresponding transformed query and key matrices become:
$\tilde{Q}=QP^{\top},\quad\tilde{K}=KP^{-1}.$ (12)
Then,
$\tilde{Q}\tilde{K}^{\top}$ $=(QP^{\top})(KP^{-1})^{\top}=QP^{\top}(P^{-1})^{\top}K^{\top}$ (13)
$=Q(P^{-1}P)^{\top}K^{\top}=QK^{\top}.$ (14)
Thus, attention logits and softmax remain unchanged. Since $V$ is untouched, the overall attention output is exactly preserved:
$\text{Attention}(\tilde{Q},\tilde{K},V)=\text{Attention}(Q,K,V).$ (15)
This proves the transformation modifies $W_{q}$ and $W_{k}$ without altering the functionality of the attention layer.
A.2 Formal Proof of Invariance
This section provides the formal proof for the invariance of our spectral fingerprints, which is a cornerstone of the method’s robustness. Our fingerprint is based on the singular value spectra of two specific matrix products, which we have selected to be strictly invariant to the corresponding transformations defined above.
Invariance of the V-O Product ($M_{ov}=W_{o}W_{v}$).
The transformation is defined on the conceptual weights as $\tilde{W}_{v}=CW_{v}$ and $\tilde{W}_{o}=W_{o}C^{-1}$. The transformed product is:
$\tilde{M}_{ov}$ $=\tilde{W}_{o}\cdot\tilde{W}_{v}$
$=(W_{o}C^{-1})\cdot(CW_{v})$
$=W_{o}\cdot(C^{-1}C)\cdot W_{v}$
$=W_{o}\cdot I\cdot W_{v}=W_{o}W_{v}=M_{ov}.$
The product matrix $M_{ov}$ is strictly invariant. Consequently, its Singular Value Decomposition (SVD), and therefore its singular value spectrum, is also invariant, regardless of the storage convention of the weights.
Invariance of the Q-K Product ($M_{q^{\top}k}=W_{q}^{\top}W_{k}$).
The transformation is $\tilde{W}_{q}=PW_{q}$ and $\tilde{W}_{k}=(P^{\top})^{-1}W_{k}$. The transformed product is:
$\tilde{M}_{q^{\top}k}$ $=\tilde{W}_{q}^{\top}\cdot\tilde{W}_{k}$
$=(W_{q}^{\top}P^{\top})\cdot(P^{\top})^{-1}W_{k}$
$=W_{q}^{\top}\cdot I\cdot W_{k}=W_{q}^{\top}W_{k}=M_{q^{\top}k}.$
A.3 Compatibility with Multi-Head Attention and Positional Encodings
The functionality-preserving transformations and their corresponding invariance proofs described previously were presented on generic linear transformation layers. In modern Transformer architectures, these transformations must be compatible with key mechanisms such as Multi-Head Attention (MHA) and Rotary Positional Embeddings (RoPE). This section details how our method ensures this compatibility.
Multi-Head Attention (MHA).
In the multi-head attention mechanism, the Query, Key, and Value matrices ($Q$, $K$, $V$) are logically partitioned into $h$ heads, with each head operating independently in a subspace of the model’s hidden dimension. To apply transformations without disrupting this structure, our transformation matrices $C$ (for V-O) and $P$ (for Q-K) are constructed as block-diagonal matrices:
$C=\begin{pmatrix}C_{1}&0&\cdots&0\\ 0&C_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&C_{h}\end{pmatrix},\quad P=\begin{pmatrix}P_{1}&0&\cdots&0\\ 0&P_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&P_{h}\end{pmatrix}$
Here, each sub-matrix $C_{i}$ and $P_{i}$ has dimensions matching that of a single head (head_dim). This design has two critical advantages:
  1. 1.
    Isolated Transformations: Each sub-matrix operates solely within the subspace of its corresponding head, never mixing information across different heads. This guarantees that the independence of the multi-head mechanism is preserved.
  2. 2.
    Preservation of Invariance: Since the overall transformation is linear and the fundamental algebraic properties (e.g., $CC^{-1}=I$) hold for block-diagonal matrices, the proofs for functionality preservation and product invariance (e.g., $\widetilde{M}_{ov}=M_{ov}$) derived in the previous section remain fully valid for the entire weight matrix.
Rotary Positional Embedding (RoPE).
RoPE injects positional information by rotating pairs of dimensions within the Query and Key vectors. For instance, it rotates dimensions $(0,1)$, then independently rotates dimensions $(2,3)$, and so on. Any transformation applied to $W_{q}$ and $W_{k}$ must respect this pairwise structure to avoid corrupting the positional information.
To this end, we impose a stronger constraint on the Q-K transformation matrix $P$, making it doubly block-diagonal. Specifically, the transformation matrix for each head, $P_{i}$, is itself a block-diagonal matrix composed of $d_{\text{head}}/2$ smaller $2\times 2$ matrices:
$P_{i}=\begin{pmatrix}P_{i,1}^{(2\times 2)}&&0\\ &\ddots&\\ 0&&P_{i,d_{\text{head}}/2}^{(2\times 2)}\end{pmatrix}$
This meticulous structure ensures that our transformation operates only within the 2D subspaces on which RoPE acts, never mixing information between them (e.g., information from dimension 1 is not leaked to dimension 2). This guarantees that RoPE can still be applied correctly to the query and key vectors after our transformation, thus strictly preserving the model’s output.
Similarly, this more granular structural constraint does not alter the fact that $P$ is an invertible matrix. Therefore, the proof of invariance for our invariant product $M_{q^{\top}k}=W_{q}^{\top}W_{k}$ remains sound.
Appendix B Methodology and Implementation Details
This appendix provides specific implementation details for the algorithms and metrics described in Section 4, as well as the hyperparameter values used in our experiments (Section 5), to ensure full reproducibility.
B.1 Core Function Definitions
Effective Rank.
The effective rank, used in Section 4.2 to determine the number of singular values to compare, is a continuous measure of a matrix’s intrinsic dimensionality, defined via the entropy of its normalized singular values. Given a singular value vector $\mathbf{s}=(s_{1},s_{2},\dots,s_{n})$, we define its effective rank as:
$\text{eff\_rank}(\mathbf{s})=\exp\left(-\sum_{j=1}^{n}\frac{s_{j}}{\sum_{i=1}^{n}s_{i}}\log\frac{s_{j}}{\sum_{i=1}^{n}s_{i}}\right)$ (16)
Distance Correlation (dCor).
Distance correlation is used to compute the ‘GhostSpec-corr‘ similarity score (Section 4.2, Equation 7). It is a measure of statistical dependence between two random vectors of arbitrary, not necessarily equal, dimension. Unlike Pearson correlation, distance correlation is zero if and only if the vectors are statistically independent, making it a robust metric for comparing the complex, non-linear trends of our layer-wise spectral mean sequences.
Formally, for two sequences $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$ with $n$ samples, the empirical distance correlation is computed as follows. First, the Euclidean distance matrices $a,b\in\mathbb{R}^{n\times n}$ are constructed, where $a_{ij}=\|x_{i}-x_{j}\|$ and $b_{ij}=\|y_{i}-y_{j}\|$. These matrices are then doubly centered:
$A_{ij}=a_{ij}-\bar{a}_{i\cdot}-\bar{a}_{\cdot j}+\bar{a}_{\cdot\cdot},$ (17)
where $\bar{a}_{i\cdot}$ is the row mean, $\bar{a}_{\cdot j}$ is the column mean, and $\bar{a}_{\cdot\cdot}$ is the grand mean of matrix $a$. The same transformation is applied to $b$ to obtain $B$. The sample distance correlation $\mathcal{R}_{n}(X,Y)$ is then defined based on the sample distance covariance $\mathcal{V}_{n}(X,Y)$ and variances $\mathcal{V}_{n}(X),\mathcal{V}_{n}(Y)$.
B.2 Hyperparameter Settings
POSA Gap Penalty ($\rho$).
The gap penalty $\rho$ for the Penalty-based Optimal Spectral Alignment (POSA) algorithm (Algorithm 1), which is central to comparing models of varying depth (Section 4.2), was set empirically. We observed that a value corresponding to the average single-layer MSE distance between moderately similar models provided a good balance between matching corresponding layers and allowing for necessary skips in pruned or expanded models. Across all experiments, a fixed value of $\rho=0.002$ was used.
Sigmoid Transformation Parameters ($\tau,k$).
The parameters for the final inverted Sigmoid transformation (Equation 6), used to convert the raw path distance from POSA into the final ‘GhostSpec-mse‘ similarity score, were determined empirically. The goal was to select values that best separate the distance distributions of related and unrelated model pairs, thereby maximizing the F1-score on a validation set. Based on this analysis, we use a fixed threshold $\tau=0.00371$ and a steepness factor $k=1000$.
Method-Specific Thresholds ($\tau^{*}$).
As described in the Evaluation Protocol (Section 5.1), we determine an optimal classification threshold $\tau^{*}$ for each fingerprinting method to ensure a fair and rigorous comparison. This threshold is found by searching for the value that maximizes the F1-score for each method on our ground-truth dataset, ensuring that every method is evaluated at its peak performance. The final, empirically determined thresholds used for all experiments are presented in Table 2.
Table 2: Optimal classification thresholds for all methods.
Method Optimal Threshold ($\tau^{*}$)
QueRE 0.18
Logits 0.83
REEF 0.83
PCS 0.12
GhostSpec-corr 0.61
GhostSpec-mse 0.85
Appendix C Detailed Experimental Setup
This appendix provides supplementary details for the experimental setup described in Section 5.1.
C.1 Model Dataset
Our comprehensive dataset consists of 63 model pairs, created using **Llama-2-7b** and **Mistral-7B** as primary base models. This dataset covers a wide range of transformations, including fine-tuning, compression, merging, and expansion, with ground-truth labels indicating whether each pair is ’related’ or ’unrelated’. A complete list of all model pairs and their corresponding modification type is provided in Table 2.
Table 3: Complete list of the model pairs used in our dataset for evaluation, supplementing Section 5.1.
Primary Model: Llama-2-7b Primary Model: Mistral-7B
Comparison Model Modification Type Comparison Model Modification Type
Llama-2-7b-chat-hf Fine-tuning OpenHermes-2.5-Mistral-7B Fine-tuning
vicuna-7b-v1.5 Fine-tuning Mistral-7B-Instruct-v0.1 Fine-tuning
Xwin-LM-7B-V0.2 Fine-tuning Nous-Hermes-2-Mistral-7B Fine-tuning
llemma_7b Fine-tuning dolphin-2.2.1-mistral-7b Fine-tuning
CodeLlama-7b-hf Fine-tuning
chinese-llama-2-7b Fine-tuning
coma-7B-v0.1 Fine-tuning
Sheared-LLaMA-1.3B Structured Pruning OpenHermes-pruned50 Pruning
Sheared-LLaMA-2.7B Structured Pruning OpenHermes-pruned2.4 Pruning
Llama-2-chat-shortgpt-30 Structured Pruning
Llama-2-pruned50-retrained Unstructured Pruning Triunvirato-7b Merging & Expansion
Llama-2-pruned70-retrained Unstructured Pruning Daredevil-7B Merging & Expansion
GBLM-Pruner-LLaMA-2-7B Unstructured Pruning BioMistral-merged-instruct Merging & Expansion
CodeLlaMa-7B-dare-ties Merging & Expansion Chunky-Lemon-Cookie-11B Merging & Expansion
llama2-7b-func-calling-slerp Merging & Expansion
FuseLLM-7B Merging & Expansion
LLaMA-Pro-8B Merging & Expansion
Camelidae-8x7B Merging & Expansion
Llama-2-7b-chat-scaled Adversarial Transforms Baichuan2-7B-Base Unrelated Models
Llama-2-7b-chat-permuted Adversarial Transforms internlm2-base-7b Unrelated Models
Mistral-7B-v0.1 Unrelated Models Qwen2.5-7B Unrelated Models
Qwen2.5-7B Unrelated Models pythia-6.9b Unrelated Models
Yi-1.5-6B Unrelated Models mpt-7b Unrelated Models
internlm2-base-7b Unrelated Models opt-6.7b Unrelated Models
pythia-6.9b Unrelated Models Phi-3-small-8k-instruct Unrelated Models
opt-6.7b Unrelated Models Yi-1.5-6B Unrelated Models
Phi-3-small-8k-instruct Unrelated Models Qwen2-7B Unrelated Models
deepseek-coder-6.7b Unrelated Models deepseek-coder-6.7b Unrelated Models
Qwen1.5-7B Unrelated Models Qwen1.5-7B Unrelated Models
C.2 Computing Infrastructure
All experiments were conducted on a server equipped with four NVIDIA GeForce RTX 4090 GPUs (24GB VRAM). The software environment was based on CUDA Version 12.2, with Python 3.10 and PyTorch 2.1.
Appendix D Additional Experiments and Visualizations
This section provides additional experimental results that empirically justify key design choices of the GhostSpec method (Section 4) and offers further visual evidence of its effectiveness.
D.1 Ablation Studies
On the Necessity of Invariant Products
To demonstrate the critical role of our invariant product construction (Section 4.1), we performed an ablation study. We compared the performance of the full GhostSpec method against a “naive” variant that computes fingerprints directly from the singular values of the raw attention weights ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$). The specific computational difference between the two methods is as follows:
  • Full GhostSpec (Invariant Method): This is the standard method proposed in our paper. For each layer $i$, it first constructs two invariant product matrices whose singular value spectra are robust to functionality-preserving transformations:
    $M_{\text{ov}}^{(i)}=W_{o}^{(i)}W_{v}^{(i)}\quad\text{and}\quad M_{q^{\top}k}^{(i)}=(W_{q}^{(i)})^{\top}W_{k}^{(i)}$ (18)
    The final fingerprint for the layer, $\mathcal{F}^{(i)}$, is then derived from the singular value vectors of these two product matrices:
    $\mathcal{F}^{(i)}=\left(\text{SVD}(M_{\text{ov}}^{(i)}),\;\text{SVD}(M_{q^{\top}k}^{(i)})\right)$ (19)
  • Naive Variant (Vulnerable Method): For the ablation study, we created a “naive” version that bypasses the invariant product construction. Instead, it directly computes the SVD of the four individual attention weight matrices for each layer $i$:
    $W_{q}^{(i)},\quad W_{k}^{(i)},\quad W_{v}^{(i)},\quad W_{o}^{(i)}$ (20)
    The fingerprint in this case is composed of the singular value vectors from these four raw matrices:
    $\begin{split}\mathcal{F}_{\text{naive}}^{(i)}=\big(&\text{SVD}(W_{q}^{(i)}),\text{SVD}(W_{k}^{(i)}),\\ &\text{SVD}(W_{v}^{(i)}),\text{SVD}(W_{o}^{(i)})\big)\end{split}$ (21)
We tested both methods on the task of identifying a transformed version of Llama-2-7b-chat-hf that had undergone both scaling and permutation attacks. The results are shown in Table LABEL:tab:ablation_results. The naive method fails completely, with similarity scores dropping significantly. In contrast, GhostSpec maintains near-perfect similarity, proving that the invariant product construction is essential for robustness against these attacks.
Table 4: Comparison between the naive direct SVD method and the invariant GhostSpec method when faced with an adversarial transformation. Scores are similarity to the original model.
Method GhostSpec-mse GhostSpec-corr
Direct SVD (Vulnerable) 0.2585 0.5684
GhostSpec (Invariant) 0.9761 1.0000
(a) Discriminative gap for Correlation-based scores.
(b) Discriminative gap for MSE-based scores.
Figure 7: Comparison of the discriminative gaps for two scoring methods. POSA achieves the largest gaps, demonstrating superior discriminative power.
On the Necessity of the POSA Algorithm.
To demonstrate the necessity of our Penalty-based Optimal Spectral Alignment (POSA) algorithm for comparing models of different depths, we conducted a targeted ablation study. The core challenge when comparing models A and B with different layer counts ($N_{A}\neq N_{B}$) is establishing a meaningful correspondence between their layer-wise feature sequences, $\mathcal{F}_{A}=(f_{A}^{(1)},\dots,f_{A}^{(N_{A})})$ and $\mathcal{F}_{B}=(f_{B}^{(1)},\dots,f_{B}^{(N_{B})})$. Our experiment was specifically designed to test this scenario.
We selected ‘Llama-2-7b‘ (32 layers) as the base model and compared it against a curated set of seven models, all of which have different layer counts, making a sophisticated alignment strategy essential. These models were categorized as:
  • Related Models: Five models derived from Llama-2 but with varying depths, including ‘LLaMA-Pro-8B‘ (40 layers), ‘Llama-2-7b-chat-hf-30-sparsity‘ (23 layers), and the ‘Sheared-LLaMA-1.3B‘ series (24 layers).
  • Unrelated Models: Two architecturally distinct models, ‘Qwen2-7B‘ (28 layers) and ‘Qwen2.5-7B‘ (28 layers).
Our experiment compared the performance of POSA against three naive baseline alignment strategies:
  • POSA (Our Method): Employs dynamic programming with a gap penalty ($\rho$) to find the optimal, potentially non-contiguous, layer alignment that minimizes global distance, as defined by the cost function:
    $\text{Cost}(P)=\sum_{(i_{k},j_{k})\in P}D_{i_{k},j_{k}}+\sum_{k>1}\rho\cdot(|j_{k}-j_{k-1}|-1)$
  • Baseline 1: Front Truncation: Compares only the first $n$ layers of each model, where $n=\min(N_{A},N_{B})$.
  • Baseline 2: Back Truncation: Compares only the last $n$ layers of each model.
  • Baseline 3: Proportional Subsampling: Downsamples the layer sequence of the larger model to match the length of the smaller one.
To quantitatively evaluate the discriminative power of each alignment strategy, we aggregated the results. For each method, we calculated the average similarity score for the “Related” group and the “Unrelated” group. The key metric for our evaluation is the Discriminative Gap ($\Delta$), which we define as the difference between these two average scores:
$\Delta=\text{AvgScore}(\text{Related})-\text{AvgScore}(\text{Unrelated})$
A larger gap indicates a superior ability to clearly and reliably distinguish genuine derivatives from unrelated models. The results of this analysis are presented in Figure LABEL:fig:posa_ablation_correlation and Figure 7(b).
As shown in the figures, our proposed POSA method consistently achieves the largest discriminative gap across both metrics. For the Correlation-based Score (Figure 7(a)), POSA achieves a gap of $\Delta=0.655$, significantly outperforming Front Truncation ($\Delta=0.322$), Back Truncation ($\Delta=0.474$), and Proportional Subsampling ($\Delta=0.457$). Similarly, for the MSE-based Score (Figure 7(b)), POSA’s gap of $\Delta=0.921$ is substantially larger than that of any other baseline.
These results empirically validate the superiority of the POSA algorithm. While naive methods can show some distinction, their separation between related and unrelated models is less pronounced and less reliable. By finding an optimal layer-wise correspondence that accounts for structural modifications, POSA maximizes the similarity signal from true derivatives while robustly rejecting unrelated models, thereby providing the most effective method for verifying model lineage when architectures differ.
D.2 Visualization of the Fingerprint Space
To visualize the global structure of the fingerprint space and the discriminative power of our method, we perform dimensionality reduction using the t-SNE algorithm. Instead of constructing explicit high-dimensional feature vectors for each model, we leverage the final similarity scores to build a comprehensive pairwise distance matrix, which directly captures the relationships between all models in our dataset. This process is executed as follows:
  1. 1.
    Pairwise Distance Matrix Construction: For a set of $N$ models in our dataset, we first compute an $N\times N$ similarity matrix $S$. Each element $S_{ij}$ in this matrix represents the similarity score between model $i$ and model $j$, calculated using either the GhostSpec-mse or GhostSpec-corr method. This similarity matrix is then converted into a distance matrix $D$, where each element is defined as:
    $D_{ij}=1-S_{ij}$
    The matrix $D$, where $D_{ii}=0$ and $D_{ij}=D_{ji}$, serves as the direct input for the t-SNE algorithm.
  2. 2.
    t-SNE Dimensionality Reduction: We use the t-SNE algorithm with the metric=’precomputed’ setting, which operates directly on our distance matrix $D$. The algorithm seeks to find a low-dimensional embedding (in our case, 2D points $Y=\{y_{1},y_{2},\dots,y_{N}\}$) that preserves the local neighborhood structure of the original data. It achieves this by minimizing the Kullback–Leibler (KL) divergence between the joint probability distribution $P$, derived from the pairwise distances in $D$, and the joint probability distribution $Q$ of the low-dimensional points in $Y$:
    $\mathrm{KL}(P\|Q)=\sum_{i\neq j}p_{ij}\log\frac{p_{ij}}{q_{ij}}$
    The resulting 2D coordinates $Y$ are then plotted, with each point representing a model.
We applied this visualization technique to a diverse dataset comprising models from the Llama family, the Mistral family, and several architecturally distinct unrelated models. The t-SNE embedding based on the GhostSpec-mse metric is shown in Figure 8. The visualization clearly demonstrates the effectiveness of our fingerprinting method. Models belonging to the Llama family form a distinct and tight cluster, while models from the Mistral family converge to a separate region in the embedded space. Critically, the unrelated models remain scattered and do not intrude upon these well-defined family clusters.
This clustering provides strong, intuitive evidence that our GhostSpec similarity scores effectively capture the underlying genealogical relationships between models. The clear separation of model families in the 2D space visually confirms that our method can robustly identify and differentiate model lineages, complementing the quantitative results presented earlier. For reproducibility, the t-SNE parameters were fixed with a perplexity of 20 and a random_state of 42.
Figure 8: t-SNE visualization of the model fingerprint space using the GhostSpec-mse distance metric. Each point represents a model, colored by its family. The clear separation between the Llama family (red), Mistral family (green), and unrelated models (blue) demonstrates that our fingerprint effectively captures genealogical relationships, causing related models to cluster together.